OpenStudy (anonymous):

-2x - 18 = 6(1 - 2x) + 4x x =

mathslover (mathslover):

\[\huge{-2x-18=6(1)-6(2x)+4x}\] \[\huge{-2x-18=6-12x+4x}\] \[\huge{-2x-18=6-8x}\] \[\huge{-2x+8x=6+18}\] \[\huge{6x=24}\] \[\huge{x=\frac{24}{6}=4}\]

OpenStudy (anonymous):

thank you !